three years ago,jim was three times as old as tom was then.In fie years,jim will be twice as old as tom.how old is each now
Answer by scott8148(6628)

j - 3 = 3(t - 3) ___ j = 3t - 6

j + 5 = 2(t + 5)

use substitution to solve